OpenStudy (anonymous):

Judith is digging for sandstone at a geological site. She has 160 meters of rope and 4 stakes to mark off a rectangular area. Which set of dimensions will create a rectangle using all the rope Judith has with her?

OpenStudy (anonymous):

65x10?

OpenStudy (anonymous):

idk is \(2\times 65+2\times 10=160\)?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

150

OpenStudy (anonymous):

then i guess that doesn't use it all up are there other choices?q

OpenStudy (anonymous):

80*80..70*10 and 16*10

OpenStudy (anonymous):

70*10

OpenStudy (anonymous):

is \[2\times 80+2\times 80=160\]?

OpenStudy (anonymous):

yeah 70 by 10 works

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw